# Complements of discriminants of real parabolic function singularities. II

### 08 Mar 2026 v5

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of eight connected domains. We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value. The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three shares of the pyramid: their intersections with a section of the pyramid are shown in Fig. 11 (right). Each share, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the saddlepoint with that marking is the highest. The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding share shown in Fig. 11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values. Clearly, all three of these domains are contractible.
@@
Proof of Lemma 50. For each $l\in L$, the orthogonal slice $U(l)$ of the discriminant stratum $\{D_{4}^{-}\}\subset\Xi_{2}$ at the point $\{F(l)\}$ is a versal deformation of the corresponding singularity of class $D_{4}^{-}$. The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant. Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph. These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials of class $\Xi_{2}$ having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
@@
Consider again the case $l\equiv x-y$. Let $f$ be a point in one of these three domains in the corresponding slice $U(l)$. We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 48, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\xi$ and $\tilde{\omega}_{1}$. For the same reasons as in the proof of Lemma 48, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex. This choice explicitly determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases. $\Box$
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
@@
Figure 12. The plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
Proof. The zero-level set of the polynomial $F(l)$ in the plane $\{l=0\}$ consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set
@@
that is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 left (in which we schematically depicted the set (64) by the union of the quadratic cone and the horizontal plane).
The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree. The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin. The set $\{F(l)=0\}$ in this surface consists of three curves that are distinguished by the conditions
(65) $\tilde{p}=0,\ \tilde{p}=\tilde{q},\ \mbox{and}\ \tilde{p}=-\tilde{q}$
and are tangent to the three lines considered in the previous paragraph. Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0. This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig. 11 left. Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig. 10 right. The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$. In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$. The three critical points with Morse index 1 occur at their intersections, see Fig. 10 (right). Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative. Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
@@
As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute. Thus, the monodromy over this circle acts in the same way on the set of three related domains in $U(l)$. We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 50. Therefore, the permuted domains belong to the set $W(l)$. $\Box$
@@
Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig. 12 right. The polynomial
@@
in some local coordinates at this point. The critical values at all other real critical points of ((66)) are positive.
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12]. The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form. Therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [22], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (66) to a polynomial whose lower weighted homogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$. A dilation of coordinates reduces it to the form (67), and technique of [3], §12, removes all higher weighted homogeneous terms of the Taylor expansion.
The set of all planes in ${\mathbb{R}}^{3}$ that are tangent to the same component of the zero-level set of (64) is connected. Therefore, the singularity will be of the same class for all other functions (66) corresponding to these planes. The positivity of other critical values follows from the same arguments as in the proof of Lemma 49. $\Box$
@@
For any linear function $l$ as in Lemma 52, the set of polynomials with $\mbox{Ind}=-3$ in a small neighborhood of the corresponding polynomial ((66)) consists of a single connected component.
Proof. By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$. Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (66). Any such slice realizes a versal deformation of the corresponding singularity. The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [26]. For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$: it is the component containing the perturbations shown in Fig. 13. According to Theorem 1.7 of [16], this component is contractible. $\Box$
@@
Let the linear function $l$ be given by ((68)). Then the set considered Lemmas 53, 54 contains a generic polynomial with the following properties:
  • it has exactly six real critical points: three with Morse index 1 and three with Morse index 2,
  • the critical values at all critical points with Morse index 1 are lower than all critical values at the critical points with Morse index 2,
@@
Proof. A polynomial with these properties can be obtained from the perturbation of the polynomial (66) that is shown in the middle part of Fig. 13 by a minor additional perturbation that properly orders the critical values. The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10]. $\Box$
Proof of Lemma 54. All these sets corresponding to different $l$ belong to the same connected component of the discriminant complement. Therefore, we can assume that $l$ is given by (68) and hence contains a polynomial satisfying the conditions listed in Proposition 55. Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions. It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
The set of all polynomials (66) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant. According to Lemma 53, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at these points. Indeed, this union is the space of a fiber bundle with contractible fibers over the space of such planes. This implies the statement of Theorem 43 for the virtual component with $\mbox{Card}=156$. $\Box$

### 16 Mar 2026 v6

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of nine connected domains: the number of domains with 0, 1, 2, 3, or 4 negative critical values is equal to 1, 1, 3, 3, or 1, respectively. We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value. The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three sectors of the pyramid: their intersections with a section of the pyramid are shown in Fig. 11 (right). Each sector, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the critical point with that marking is the highest. The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding sector shown in Fig. 11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values. Clearly, all three of these domains are contractible. The set of polynomials (64) belongs to the boundaries of all these domains.
Thus, we have the following statement.
Proposition 53.
There are natural one-to-one correspondences between the following three-element sets:
  1. (1)
    The set of connected components of the space of non-discriminant polynomials ((62)) having a minimum point and three critical points with Morse index 1, only one of whose critical values is positive;
  2. (2)
    The set of critical points with Morse index 1 of an arbitrary polynomial ((62)) having a minimum point;
  3. (3)
    The set of pairs of lines of the zero-level set of the polynomial $X^{3}-3XY^{2}+Z^{2}$ in the plane $\{Z=0\}.$
Proof. The sets mentioned in item 2 for all polynomials (62) having a minimum point are in a standard one-to-one correspondence with each other. This correspondence is defined by the continuity within the contractible set of all such polynomials. Thus, it is sufficient to define the correspondence $1\leftrightarrow 2$ assuming that the polynomial mentioned in item 2 belongs to this component. In this case, we associate with the connected component the critical point with a positive critical value. Also, it is sufficient to define the correspondence $2\leftrightarrow 3$ assuming that the polynomial (62) satisfies the conditions (64). In this case, the zero-level set of such a polynomial in the plane $\{Z=0\}$ looks as in Fig. 10 (right). The desired correspondence associates with each critical point the pair of lines of Fig. 10 (left) that are parallel to the lines in Fig. 10 (right) passing through this point. $\Box$
@@
Proof of Lemma 52. For each $l\in L$, the orthogonal slice $U(l)$ of the discriminant stratum $\{D_{4}^{-}\}\subset\Xi_{2}$ at the point $\{F(l)\}$ is a versal deformation of the corresponding singularity of class $D_{4}^{-}$. The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant. This map defines a bijection between the union of three domains mentioned in item 1 of Proposition 53 and the union $V(l)\cup W(l)$, see Definition 12.
@@
Consider again the case $l\equiv\frac{x-y}{\sqrt{2}}$. Let $f$ be a point of the set $V(l)\cup W(l)$ it the corresponding slice $U(l)$. We assume that this point $f$ is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 49, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\xi$ and $\tilde{\omega}_{1}$. For the same reasons as in the proof of Lemma 49, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the component containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like rightmost vertex. This choice uniquely determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Indeed, the information on this polynomial known to us determines all the elements of the virtual function up to the surgeries not crossing the discriminant: all these surgeries consist in collisions of the critical values at distant critical points. Starting from an arbitrary element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the rightmost vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other two cases. $\Box$
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the orthogonal four-dimensional slices $U(l)$ of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
@@
Figure 12. Plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
Proof. For any $l\in L$, the subset in ${\mathbb{R}}^{3}$ defined by equations
(65) $l=0,F(l)=0$
consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set where
@@
that is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 (left) (in which we schematically depicted the set (66) by the union of the quadratic cone and the horizontal plane).
Proposition 53 reduces the monodromy action on the set of components of the space $V(l)\cup W(l)\subset U(l)$ to the action on the set of pairs of these lines. For any $\varkappa=(\varkappa_{1},\dots,\varkappa_{4})$, denote by $g_{\varkappa}$ the corresponding polynomial (62). According to the definition of versal deformations, there is a neighborhood $\tilde{U}$ in the space ${\mathbb{R}}^{4}$ of parameters $\varkappa$, a diffeomorphism $\theta:\tilde{U}\to U(l)$, and a family of local diffeomorphisms $\Gamma_{\varkappa}$ from a neighborhood of the origin in the space of arguments $(X,Y,Z)$ of polynomials (62) to a neighborhood of the origin in the space of arguments $(x,y,z)$ of polynomials of class $\Xi_{2}$, such that the diffeomorphisms $\Gamma_{\varkappa}$ depend smoothly on the parameters $\varkappa\in\tilde{U}$, and for any $\varkappa\in\tilde{U}$ we have
(67) $g_{\varkappa}(X,Y,Z)\equiv f_{\theta(\varkappa)}\circ\Gamma_{\varkappa}(X,Y,Z).$
The diffeomorphism $\Gamma_{0}$ can be considered as a change of local coordinates that reduces the polynomial $F(l)$ to the normal form $X^{3}-3XY^{2}+Z^{2}$. The kernel of the second differential of a polynomial is independent of the coordinates, therefore, this diffeomorphism maps the plane $\{Z=0\}$ to a surface that is tangent to the plane $\{l=0\}$. It also maps the three lines distinguished by the equation $g_{0}=0$ in the plane $\{Z=0\}$ to three curves that are tangent to three lines forming the set (65). The map $\theta$ defines a bijection between the set of polynomials $g_{\varkappa}\subset\tilde{U}$ having a minimum point and $\mbox{Ind}=-1$, on the one hand, and the union of the subsets $V(l)$ and $W(l)$ of the slice $U(l)$, on the other hand. According to Proposition 53, we obtain a one-to-one correspondence between the pairs of lines forming the set (65) and the set of connected components of the space $V(l)\cup W(l)$.
This correspondence does not depend on the choice of maps $\Gamma_{\varkappa}$ and $\theta$ reducing the versal deformations (62) and $U(l)$ to one another. Indeed, let $\bar{f}\in U(l)$ be an arbitrary polynomial very close to $F(l)$, which has a minimum point and three critical points $a,b,$ and $c$ with Morse index 1 and critical value 0. (Such points form a single curve in $U(l)$ which is the image of the curve (64) under the map $\theta$.) To each of the points $a,b,c$, a connected component of the set $V(l)\cup W(l)$ corresponds: it contains all polynomials neighboring to $\bar{f}$ and such that the critical values of their critical points neighboring to the selected critical point of $\bar{f}$ is positive, and the critical values at the other two critical points with Morse index 1 are negative. Also, a pair of lines forming the set (65) corresponds to each point $a,b,$ and $c$. These two lines are nearly parallel to the two edges of the triangle $\triangle_{abc}\subset{\mathbb{R}}^{3}$ that contain this point. The composition of these two correspondences defines an invariant description of the correspondence between the components of $V(l)\cup W(l)$ and the pairs of lines in set (65).
Thus, the monodromy action of the circle $L/{{\mathbb{Z}}_{2}}$ on the set of pairs of lines in (65) coincides with its action on the set of corresponding connected components of the set $V(l)\cup W(l)$.
@@
As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these lines in $\ker l$ move in such a way that one pair of them returns to itself, while the other two pairs permute. Thus, the monodromy over this circle acts in the same way on the set of three related domains in $U(l)$. We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 52. Therefore, the permuted domains belong to the set $W(l)$. $\Box$
@@
Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary non-zero linear function, whose kernel is tangent to the component of the surface ((66)) that is homeomorphic to the quadratic cone, see Fig. 12 ((right)). The polynomial
@@
in some local coordinates at this point. The critical values of the polynomial ((68)) at all its other real critical points are positive.
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ or $E_{6}$, see [21], [12]. The restriction of the polynomial ((68)) to the plane $\ker l$ is a degenerate cubic form. Therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [23], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (68) to a polynomial whose lower weighted homogeneous part is $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$. A dilation of coordinates reduces it to the form (69), and technique of [3], §12, removes all higher weighted homogeneous terms of the Taylor expansion.
The set of all polynomials (68) such that the plane $\ker l$ is tangent to the same component of the zero-level set of (66) is connected. Therefore, the singularity will be of the same class for all other functions (68). The positivity of other critical values follows from the same arguments as in the proof of Lemma 50. $\Box$
@@
For any linear function $l$ as in Lemma 55, the set of polynomials with $\mbox{Ind}=-3$ has exactly one local connected component near the corresponding polynomial ((68)) in the space $\Xi_{2}$.
Proof. By the versality of the deformation (42), the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$. Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (68). Any such slice realizes a versal deformation of the corresponding singularity. The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [27]. For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$.
@@
Specifically, in the case of any singularity in two variables that is equivalent to $\xi^{2}\eta-\eta^{4}$ it is the component containing the perturbations whose zero-level sets are as shown in Fig. 13 (right). $\Box$
@@
Let the linear function $l$ be given by ((70)). Then the set considered in Lemmas 56 and 57 contains a generic polynomial with the following properties:
  • it has exactly six real critical points: three with Morse index 1 and negative critical values, and three with Morse index 2 and positive critical values;
@@
Proof. According to the Gusein-Zade—A’Campo method [10], the intersection matrix of five cycles vanishing in the splitting of $D_{5}^{-}$ singularity shown in Fig. 13 is given by the Coxeter–Dynkin diagram , where white (respectively, black) circles correspond to the critical points with Morse index 1 and negative critical value (respectively, with Morse index 2 and positive critical value). Using the Lyashko–Looijenga map, we can achieve that their critical values are ordered as follows: 53421, which corresponds to the matrix (71). All these perturbations can be performed in an arbitrarily small neighborhood of the polynomial (68), so that they will not create real critical points in addition to six critical points of this polynomial. $\Box$
Proof of Lemma 57. All these sets corresponding to different planes $\ker l$ belong to the same connected component of the discriminant complement. Therefore, we can assume that $l$ is given by (70), and any neighborhood of the polynomial (68) contains a polynomial satisfying the conditions listed in Proposition 58. Our program, upon a special request, listed all virtual functions of $P_{8}^{2}$ class satisfying these conditions. There are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
The set of all polynomials (68) with $\ker l$ tangent to the cone is connected and $S(3)$-invariant. According to Lemma 56, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at the points corresponding to these polynomials. Indeed, this union is the space of a fiber bundle with connected base and fibers. This implies the statement of Theorem 44 for the virtual component with $\mbox{Card}=156$. $\Box$
Conjecture 59.
The connected components with Card equal to 1216, 1318, 844, or 156 have non-trivial first homology groups.